OpenStudy (anonymous):

Algebra 2 help please

OpenStudy (anonymous):

Mrs. Jones Algebra 2 class scored very well on yesterday’s quiz. With one exception, everyone received an A. Within how many standard deviations of the mean do all the quiz grades fall? 91, 92, 94, 88, 96, 99, 91, 93, 94, 97, 95, 97 (1 point) 2 1 3 4

OpenStudy (anonymous):

calculate the mean first

OpenStudy (anonymous):

93.9 = 94

OpenStudy (anonymous):

now minus it from each value and square it

OpenStudy (anonymous):

like (91-94)^2 = (-3)^2 =9

OpenStudy (anonymous):

so i do that for all of them?

OpenStudy (anonymous):

yes for all

OpenStudy (anonymous):

and calculate their mean that's your standard deviation

OpenStudy (anonymous):

9,4,0,36,4,25,9,1,0,9,1,9

OpenStudy (anonymous):

98.7=99

OpenStudy (anonymous):

shouldn't it be 8?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

9,4,0,36,4,25,9,1,0,9,1,9 calculate mean of these

OpenStudy (anonymous):

i did and i got 98.7=99

OpenStudy (anonymous):

add them and divide by 12 @electrokid how do you find how many standard deviations are there

OpenStudy (anonymous):

and square-root

OpenStudy (anonymous):

sd dev = sqrt of ( mean of ( squares of difference of each value from its mean

OpenStudy (anonymous):

oh yeah sorry

OpenStudy (anonymous):

and how to find the number of standard deviations? @electrokid

OpenStudy (anonymous):

@aajugdar what do you mean number of std dev?

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

Within how many standard deviations of the mean do all the quiz grades fall?

OpenStudy (anonymous):

@vivianrush you did correct... looks like all that is left is a square-root

OpenStudy (anonymous):

@aajugdar oh. k max sample value / std dev

OpenStudy (anonymous):

and round it up

OpenStudy (anonymous):

@vivianrush \[\sqrt{(9+4+0+36+4+25+9+1+0+9+1+9)/12}\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

0.86

OpenStudy (anonymous):

~_~ its not 0.86

OpenStudy (anonymous):

its 8.91 take it as 9 take root of it ans =3

OpenStudy (anonymous):

so @electrokid 99/3 = 33 what after this

OpenStudy (anonymous):

yes a little more than 3. and the ans = max diff from mean / std dev

OpenStudy (anonymous):

no.. the question asks for $$\sigma_{mean}=\frac{\sigma}{\sqrt{N}}$$ sigma = what you found above N = 12

OpenStudy (anonymous):

so its 3/sqrt12

OpenStudy (anonymous):

its less than 1 :/

OpenStudy (anonymous):

ok,almost 1 considering SD is more than 3

OpenStudy (anonymous):

mean / std dev

OpenStudy (anonymous):

yes that should be it

OpenStudy (anonymous):

so answer will be 1

OpenStudy (anonymous):

im still confused,but thank you guys

OpenStudy (anonymous):

@vivianrush @aajugdar (99-93.9)/3.1 = 5.1/3.1 = 2 (within two standard deviations from mean)

OpenStudy (anonymous):

haha xD @vivianrush calculate mean of samples then minus it from samples and square then you got these 9,4,0,36,4,25,9,1,0,9,1,9 now calculate their mean and calculate its root you will get standard deviation and then use the formula that electro gave

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):